Parth (parthkohli):

\[\sqrt{2\sqrt{2\sqrt{2\sqrt{2}}}} \]

OpenStudy (anonymous):

what to do with Parth?

Parth (parthkohli):

Simplify :P

Parth (parthkohli):

How?

OpenStudy (lgbasallote):

that doesnt go on forever in a loop?

Parth (parthkohli):

Nope...

OpenStudy (anonymous):

\[\huge 2^{\frac{1}{2}}\times2^{\frac{1}{4}}\times2^{\frac{1}{8}}\times2^{\frac{1}{16}}\]

Parth (parthkohli):

Wow, is that really it? :P

OpenStudy (anonymous):

yeah

OpenStudy (unklerhaukus):

\[\sqrt{2\sqrt{2\sqrt{2\sqrt{2}}}}\] \[=\sqrt{2\sqrt{2\sqrt{8}}}\]\[=\sqrt{2\sqrt{32}}\]\[=\]

Parth (parthkohli):

Oh... that's a nice approach!

Parth (parthkohli):

\[\sqrt{2\sqrt{32} } \implies \sqrt{128} \]

OpenStudy (anonymous):

man wait thats wrong...

OpenStudy (unklerhaukus):

yeah my method is wrong

OpenStudy (unklerhaukus):

\[\sqrt{2\sqrt{2\sqrt{2\sqrt{2}}}}\] \[=\sqrt{2\sqrt{2\sqrt{\sqrt{8}}}}\]\[=\sqrt{2\sqrt{\sqrt{\sqrt{16\times8}}}}\]\[=\sqrt{\sqrt{\sqrt{\sqrt{256\times16\times8}}}}\]